Which of the following ratios are equivalent to 12/20 ?
A.6/10
B.5:3
C.120 to 200
D.60/80

Answer by jim_thompson5910(35256)

12/20 = 6/10 (divide both by 2)

12/20 is NOT equivalent to 5:3 since the first number of 12/20 is smaller

12/20 = 120/200 (multiply both by 10)

12/20 = 60/100, so 12/20 is NOT equivalent to 60/80